OpenStudy (anonymous):

2(n - 3) + 5 = 3(n - 1)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 2n - 6 + 5 = 3n - 3 }\) \(\Large \color{purple}{\rightarrow 2n - 1 = 3n - 3 }\) \(\Large \color{purple}{\rightarrow 2n = 3n - 2 }\) \(\Large \color{purple}{\rightarrow -n = -2 }\) \(\Large \color{purple}{\rightarrow n = 2 }\)

OpenStudy (callisto):

2(n - 3) + 5 = 3(n - 1) 2n - 2(3) +5 = 3n - 3(1) 2n -6 +5 = 3n -3 2n -1 =3n -3 Can you continue?

OpenStudy (anonymous):

2n - 6 + 5 = 3n - 3 -3n -1n-6+5 = -3 -1n-1 = -3 -1n=-2 n=2

OpenStudy (anonymous):

thanks!